A petite young woman distributes her 510 N weight equally over the heels of her high-heeled shoes. Each heel has an area of 0.850 cm2 .

What pressure is exerted on the floor by each heel?

With the same pressure, how much weight could be supported by two flat-bottomed sandals, each of area 190 cm2 ?
Answer #1

Pressure exerted on the floor by each heel,

P = (W/2) / A = (510 / 2) / (0.850 * 10-4) = 3 * 106 Pa

Weight supported by flat bottomed sandals,

W = 2 * PA = 2 * (3 * 106) * (190 * 10-4) = 114000 N = 114 kN